OpenStudy (anonymous):

Trig/ Law of Sine/ SSA / How does this have two angles? Drawing coming up and details

OpenStudy (anonymous):

 a=7 b=5 A=70 I know how we get B, which is sin B 5sin 70 / 7 which is 42.20 but what or how to we get the second B ?

OpenStudy (anonymous):

The line in the center is h

geerky42 (geerky42):

What do you mean by "second B"?

OpenStudy (anonymous):

\( \frac{5\sin 70}{7} = B_1 = 42.20\) Now they stat there is a \( B_2 = 137.80\) but I do not know where that is coming from

geerky42 (geerky42):

So you are looking for a reason why \(B_2 = 137.80\)? Can you label \(B_2\)? Where is it in triangle you sketched?

OpenStudy (anonymous):

that is not labeled.

OpenStudy (anonymous):

They stat there are two angles for B 0 < B < 180

OpenStudy (anonymous):

They stat there are two angles B 0 < B < 180, for which sin B = \(\frac{5\sin 70}{7} \)= 42.20

OpenStudy (anonymous):

First angle is 42.20 how to find the other?

OpenStudy (anonymous):

Ah I got it HAHAHA

geerky42 (geerky42):

That's good then. We don't know what second angle is referred to...

OpenStudy (anonymous):

To find the other angle we do 180 - 42.20

OpenStudy (anonymous):

That is how you get the second angle when doing SSA